OpenStudy (meilendurcer):

1/4 + 3/8= A. 1/3 B. 2/3 C. 3/4 D. 3/8 my answer is A is that correct?

OpenStudy (shane_b):

Something's wrong here...none of those answers match. \[\frac{1}{4}+\frac{3}{8}==>\frac{2}{8}+\frac{3}{8}=\frac{5}{8}\]

OpenStudy (meilendurcer):

i know right

OpenStudy (shane_b):

Maybe D is a typo and should read 5/8s....who knows

OpenStudy (meilendurcer):

yea ok thanks :)

OpenStudy (shane_b):

np :)